# 11footnotetext: Equal Contributions

### v4

@@ Abstract
We address key challenges in Dataset Aggregation (DAgger) for real-world contact-rich manipulation: how to collect informative human correction data and how to effectively update policies with this new data. We introduce Compliant Residual DAgger (CR-DAgger), which contains two novel components: 1) a Compliant Intervention Interface that leverages compliance control, allowing humans to provide gentle, accurate delta action corrections without interrupting the ongoing robot policy execution; and 2) a Compliant Residual Policy formulation that learns from human corrections while incorporating force feedback and force control. Our system significantly enhances performance on precise contact-rich manipulation tasks using minimal correction data, improving base policy success rates by over 60% on two challenging tasks (book flipping and belt assembly) while outperforming both retraining-from-scratch and finetuning approaches. Through extensive real-world experiments, we provide practical guidance for implementing effective DAgger in real-world robot learning tasks. Result videos are available at: https://compliant-residual-dagger.github.io/
@@ 1 Introduction
Learning from human demonstrations has seen many recent successes in real-world robotic tasks [chi2024universal, hou2024adaptive, wu2025robocopilot, xu2025robopanoptes, xiong2025vision]. However, to obtain a successful policy, human demonstrators often have to repeatedly deploy a policy and observe its failure cases, then collect more data to update the policy until it succeeds. This process is broadly referred to as Dataset Aggregation (DAgger) [ross2011reduction, kelly2019hg]. However, doing DAgger effectively for real-world robotic problems still faces the following challenges:
How to collect informative human correction data? DAgger is most effective when the correction data is within the original policy’s induced state-action distribution [ross2011reduction]. In practice, the common approach is either (1) collecting offline demonstrations that cover the policy’s typical failure scenarios [chi2023diffusion], or (2) human taking over robot control during policy deployment [spencer2020learning, mandlekar2020human]. However, in both cases, the human demonstrator has no access to the original policy’s behavior and may deviate excessively from it. Human taking over additionally introduces force discontinuity when they do not instantly reproduce the exact same robot force. This is partially due to the lack of effective correction interfaces that support precise and instantaneous intervention.
How to effectively update the policy with new data? Prior methods for improving a pretrained policy with additional data include (1) retraining the policy from scratch with the aggregated dataset [kelly2019hg], which can be computationally expensive; (2) finetuning the policy with only the additional data [wu2025robocopilot, he2025uncertainty, chen2025conrft], which is sensitive to the quality of the new data [yuan2024policy], and (3) training a residual policy separately on top of the pretrained policy, which is typically done with Reinforcement Learning  [ankile2024imitation, yuan2024policy] or Imitation Learning [bharadhwaj2024track2act], both require a large number of samples.
@@
Together, our system significantly improves the success rate of precise contact-rich robot manipulation tasks using a small amount of additional data. We demonstrate the efficacy of our method on two challenging tasks involving long horizons and sequences of contacts: book flipping and belt assembly. We improve over the base policy success rate by over 60%, while also outperforming retrain-from-scratch and finetuning under the same data budgets. In summary, our contributions are:
@@ 2 Related Work
Human-in-the-Loop Corrections for Robot Policy Learning. The original DAgger work [ross2011reduction] requires the demonstrator to directly label actions generated by the policy. In robotics, a practical variation is to let the human take over the robot control and provide correct action directly [kelly2019hg]. Such human correction motion can be recorded with spacemouse [chen2025conrft, liu2022robot], joystick [spencer2020learning], smartphone [mandlekar2020human], or arm-based teleoperation system [hoque2021thriftydagger, hoque2021lazydagger, wu2025robocopilot]. We instead proposes a novel kinesthetic teaching system with compliance controller that allows the demonstrator to apply delta corrections while the robot policy is still running, and additionally records force feedback. Our results show that both the delta correction data and the force data are crucial to the success of the learned policy.
@@
Improving Pretrained Robot Policies with New Data. The most direct approach to improving a pretrained policy with new correction data is to retrain the policy on the aggregated dataset, combining prior demonstrations with new feedback [mandlekar2020human, spencer2020learning]. Alternatively, Reinforcement Learning (RL) offers a framework to incorporate both offline and online data, either by warm-starting replay buffers [luo2024precise, ball2023efficient] or by using offline data to guide online fine-tuning [yin2025rapidly, xu2022aspire]. When policies are trained on large-scale human demonstration datasets [black2024pi_0, team2024octo, kim2024openvla, o2024open, khazatsky2024droid, xu2023xskill, xu2024flow], retraining becomes impractical, especially when the original data is inaccessible. In such cases, fine-tuning with only the new data is a common solution, using either imitation learning [liu2022robot, he2025uncertainty, wu2025robocopilot] or RL [mark2024policy, chen2025conrft]. Another line of work introduces an additional residual model on top of the original policy. These residual policies can be trained with RL in simulation [yuan2024policy, ankile2024imitation, haldar2023watch], but suffers from sim-to-real challenges. Training residual policy in the real world usually requires a large number of samples [bharadhwaj2024track2act, johannink2019residual], intermediate scene representation [guzey2024bridging], or consistent visual observations between training and testing [guzey2024see, haldar2023teach], making the approach hard to adopt in practice. In this work, we introduce a practical data collection system and an efficient residual policy learning algorithm for long-horizon, contact-rich manipulation tasks. Our approach requires only a small amount of real-world correction data and supports integration of additional sensory modalities not present in the original model, leading to improved policy performance.
@@ 3.1 Compliant Intervention Interface
Correction data is collected by human demonstrators to rectify policy failures. Unlike initial demonstrations that establish baseline behaviors, correction data specifically targets failure modes observed during policy deployment. Correction data is most effective when it corrects failures in policy-induced state distributions [ross2011reduction]. The interface through which these corrections are collected significantly impacts the quality of correction data, which should be intuitive for demonstrators, capture critical corrective information at precise moments of failures, and facilitates collecting data close to the base policy state-action distribution.
@@
  • Indirect correction brings errors. Correction is commonly implemented via teleoperation interfaces such as spacemouse or joysticks [he2025uncertainty, chen2025conrft]. With spatial mismatch and teleoperation latency, it is hard for the demonstrator to instantly provide accurate corrections upon intervention starts without going through a short adjustment period.
@@
  • Correction interface with compliance control. In order to apply delta correction over a running policy, we provide a compliant interface that allows humans to safely intervene and apply force to the robot to affect its behaviors at any time, as shown in Fig. 2. We design a kinesthetic correction hardware setup with a detachable handle for human to hold when correcting, and allows easy tool-swapping for different tasks. We run a compliance controller (specifically admittance control) in the background to respond to both contact forces and human correction forces, allowing the human to influence but not completely override the policy execution. The admittance controller uses a constant stiffness $\sim$$1000\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$ to allow easy human intervention and ensure accurate tracking.
@@ 3.2 Compliant Residual Policy
Compliant residual policy formulation. Our policy directly learns corrective behavior from the human delta correction data, as shown in Fig. 3. It takes as input the same visual and proprioceptive feedback as the base policy but with a shorter horizon. It also takes in an extra force modality, which is available using our Compliant Intervention Interface. The policy outputs five frames of actions at a time, corresponding to $0.1\text{\,}\mathrm{s}$ of execution time when running at $50\text{\,}\mathrm{H}\mathrm{z}$. The action space is 15-dimensional: the first nine dimensions represent the SE3 delta pose from the base policy action to the robot pose command [chi2023diffusion], while the later six dimensions represent the expected wrench (force and torque) the robot should feel from external contacts. Both the robot pose command and the expected wrench are sent to a standard admittance controller for execution with compliance.
The residual policy directly uses the base policy’s frozen image encoder [radford2021learning, amir2021deep, xu2023jacobinerf] to extract an image embedding, a temporal convolution network [van2016wavenet] to encode the force vectors, followed by fully-connected layers to decode actions.
@@
Training strategy. In prior work, a residual policy is trained either in simulation with RL [ankile2024imitation, yuan2024policy] to give it sufficient coverage of the state distribution, or in the real world with pre-collected behavior cloning data [luo2024precise]. In this work, we train the Compliant Residual Policy completely on the small amount of new real-world correction data with the following strategies:
@@
  • Prioritize correction data over no-correction (zero residual action) data. Similar to [liu2022robot], we alter the data sample frequency during training based on whether they have human correction or not. Specifically, since the moment of correction start indicates where the current policy performs badly followed by immediate action to fix it, we sample data more frequently for a short period immediately after correction starts. Our real-world ablations (§ 4.5) demonstrate that our training strategies improve the quality of the residual policy and the overall success rate.
@@ 4.1 Contact-Rich Manipulation Tasks
This task is challenging as it involves rich use of physical contacts and forceful strategies [hou2020manipulation]. A position-only strategy always fails immediately by triggering large forces, so we execute all policies through the same admittance controller. The task success requires high precision in both motion and force to accurately align the fingers with the gap upon insertion, and to provide enough force to rotate heavy books and make the books stand firmly.
@@
Belt Assembly: As shown in Fig. 5 (a), this task is to assemble a thin belt onto two pulleys, which is part of the NIST board assembly challenge [kimble2020benchmarking]. Starting with the belt grasped by the gripper, the robot needs to first thread the belt over the small pulley, next move down while stretching the belt to thread its other side on the big pulley, then rotate $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ around the big pulley to tighten the belt, and finally pull up to release the belt from the gripper. The base policy is trained with 405 demonstrations.
@@ 4.2 Base Policy and its Failure Modes
We trained a diffusion policy [chi2023diffusion] that takes in past images from a wrist-mounted camera and robot proprioception observations, and predicts a future position-based action trajectory. To isolate the contribution of force inputs versus human corrections, we trained diffusion policies with and without force inputs as baselines for the belt assembly task.
@@ 4.3 Comparisons
  • Take-over-Correction (HG-DAgger) [kelly2019hg] is another common correction strategy where human demonstrators monitor policy execution and take complete control when failures are anticipated. We implement Take-over-Correction on our Compliant Intervention Interface by cleaning up command buffer to the compliance controller and switching stiffness to zero upon correction starts, so the robot policy does not affect the robot during correction. However, as explained in § 3.1, take-over correction introduces an abrupt transition around control authority switching, which may cause distributional discontinuities in the training data.
@@
  • Fintune Policy with KL Regularization: A recent method [fan2023dpok] that stabilizes finetuning training by encouraging the predicted action to be close to the training data distribution.

### v5

# Compliant Residual DAgger: Improving Real-World Contact-Rich Manipulation with Human Corrections

@@ Abstract
We address key challenges in Dataset Aggregation (DAgger) for real-world contact-rich manipulation: how to collect informative human correction data and how to effectively update policies with this new data. We introduce Compliant Residual DAgger (CR-DAgger), which contains two novel components: 1) a Compliant Intervention Interface that leverages compliance control, allowing humans to provide gentle, accurate delta action corrections without interrupting the ongoing robot policy execution; and 2) a Compliant Residual Policy formulation that learns from human corrections while incorporating force feedback and force control. Our system significantly enhances performance on precise contact-rich manipulation tasks using minimal correction data, improving base policy success rates by 64% on four challenging tasks (book flipping, belt assembly, cable routing, and gear insertion) while outperforming both retraining-from-scratch and finetuning approaches. Through extensive real-world experiments, we provide practical guidance for implementing effective DAgger in real-world robot learning tasks. Result videos are available at: https://compliant-residual-dagger.github.io/
keywords:
DAgger, Imitation Learning, Manipulation, Compliance Control
@@ 1 Introduction
Learning from human demonstrations has seen many recent successes in real-world robotic tasks chi2024universal; hou2024adaptive; wu2025robocopilot; xu2025robopanoptes; xiong2025vision. However, to obtain a successful policy, human demonstrators often have to repeatedly deploy a policy and observe its failure cases, then collect more data to update the policy until it succeeds. This process is broadly referred to as Dataset Aggregation (DAgger) in ross2011reduction; kelly2019hg. However, doing DAgger effectively for real-world robotic problems still faces the following challenges:
How to collect informative human correction data? ross2011reduction shows that DAgger is most effective when the correction data is within the original policy’s induced state-action distribution. In practice, the common approach is either (1) collecting offline demonstrations that cover the policy’s typical failure scenarios chi2023diffusion, or (2) human taking over robot control during policy deployment spencer2020learning; mandlekar2020human. However, in both cases, the human demonstrator has no access to the original policy’s behavior and may deviate excessively from it. Human taking over additionally introduces force discontinuity when they do not instantly reproduce the exact same robot force. This is partially due to the lack of effective correction interfaces that support precise and instantaneous intervention.
How to effectively update the policy with new data? Prior methods for improving a pretrained policy with additional data include (1) retraining the policy from scratch with the aggregated dataset kelly2019hg, which can be computationally expensive; (2) finetuning the policy with only the additional data wu2025robocopilot; he2025uncertainty; chen2025conrft, which is sensitive to the quality of the new data yuan2024policy, and (3) training a residual policy separately on top of the pretrained policy, which is typically done with Reinforcement Learning  ankile2024imitation; yuan2024policy or Imitation Learning bharadhwaj2024track2act, both require a large number of samples.
@@
Together, our system significantly improves the success rate of precise contact-rich robot manipulation tasks using a small amount of additional data. We demonstrate the efficacy of our method on four challenging tasks involving long horizons and sequences of contacts: book flipping, belt assembly, cable routing, and gear insertion. We improve over the base policy success rate by 64%, while also outperforming retrain-from-scratch and finetuning under the same data budgets. In summary, our contributions are:
@@
This work is an extended version of the conference paper CRDagger2025. We expand the content of this paper in the following ways:
  • Enrich the experimental results with two additional contact-rich tasks as detailed in Sec. 4: a peg-in-hole style gear insertion task, which requires sub millimeter accuracy and tests our system for small-magnitude correction motions; and a cable routing task, which contains large variations in item physical properties (cable stiffness).
  • Identify and explain the intention-misinterpretation phenomenon in correction data collection in Sec. 3.1, caused by robot tracking errors. We provide a solution to the problem by reducing tracking error both spatially and temporally and include additional analysis on the effect of tracking error in Sec. 4.5.
  • Analyze the quality improvement of human correction data from our proposed system comparing with traditional manipulation data collection method in Sec. 4.4.
@@ 2 Related Work
Human-in-the-Loop Corrections for Robot Policy Learning. The original DAgger work ross2011reduction requires the demonstrator to directly label actions generated by the policy. In robotics, a practical variation is to let the human take over the robot control and provide correct action directly kelly2019hg. Such human correction motion can be recorded with spacemouse chen2025conrft; liu2022robot, joystick spencer2020learning, smartphone mandlekar2020human, or arm-based teleoperation system hoque2021thriftydagger; hoque2021lazydagger; wu2025robocopilot. We instead proposes a novel kinesthetic teaching system with compliance controller that allows the demonstrator to apply delta corrections while the robot policy is still running, and additionally records force feedback. Our results show that both the delta correction data and the force data are crucial to the success of the learned policy.
@@
Improving Pretrained Robot Policies with New Data. The most direct approach to improving a pretrained policy with new correction data is to retrain the policy on the aggregated dataset, combining prior demonstrations with new feedback mandlekar2020human; spencer2020learning. Alternatively, Reinforcement Learning (RL) offers a framework to incorporate both offline and online data, either by warm-starting replay buffers luo2024precise; ball2023efficient or by using offline data to guide online fine-tuning yin2025rapidly; xu2022aspire. When policies are trained on large-scale human demonstration datasets black2024pi_0; team2024octo; kim2024openvla; o2024open; khazatsky2024droid; xu2023xskill; xu2024flow, retraining becomes impractical, especially when the original data is inaccessible. In such cases, fine-tuning with only the new data is a common solution, using either imitation learning liu2022robot; he2025uncertainty; wu2025robocopilot or RL mark2024policy; chen2025conrft. Another line of work introduces an additional residual model on top of the original policy. These residual policies can be trained with RL in simulation yuan2024policy; ankile2024imitation; haldar2023watch, but suffers from sim-to-real challenges. Training residual policy in the real world usually requires a large number of samples bharadhwaj2024track2act; johannink2019residual, intermediate scene representation guzey2024bridging, or consistent visual observations between training and testing guzey2024see; haldar2023teach, making the approach hard to adopt in practice. In this work, we introduce a practical data collection system and an efficient residual policy learning algorithm for long-horizon, contact-rich manipulation tasks. Our approach requires only a small amount of real-world correction data and supports integration of additional sensory modalities not present in the original model, leading to improved policy performance.
Compliance and Compliance Control. Compliance is a motor property that describes how motion responds to force. For example, traditional industrial robots typically have position control with very low compliance to achieve precise tracking under heavy load. On the contrary, high compliance lets a robot retreat when it experiences external forces. A high compliance or a variable compliance profile mason2007compliance may be preferred for tasks involving interactions with the environment, since they can increase execution robustness or avoid huge contact forces during the interactions hou2019robust; hou2020manipulation.
Compliance control refers to feedback control techniques that give a robot high compliance or spatial-temporally varying compliance profiles. Compliance control can be implemented by Impedance Control hogan1984impedance if the robot is back-drivable, or by Admittance Control with external force sensors if the robot cannot be back-driven, such as most industrial robots. A more through review of compliance controller implementations can be found in lynch2017modern and villani2016force.
@@ 3.1 Compliant Intervention Interface
Correction data is collected by human demonstrators to rectify policy failures. Unlike initial demonstrations that establish baseline behaviors, correction data specifically targets failure modes observed during policy deployment. Correction data is most effective when it corrects failures in policy-induced state distributions ross2011reduction. The interface through which these corrections are collected significantly impacts the quality of correction data, which should be intuitive for demonstrators, capture critical corrective information at precise moments of failures, and facilitates collecting data close to the base policy state-action distribution.
@@
  • Indirect correction brings errors. Correction is commonly implemented via teleoperation interfaces such as spacemouse or joysticks he2025uncertainty; chen2025conrft. With spatial mismatch and teleoperation latency, it is hard for the demonstrator to instantly provide accurate corrections upon intervention starts without going through a short adjustment period.
@@
  • Correction interface with compliance control. In order to apply delta correction over a running policy, we provide a compliant interface that allows humans to safely intervene and apply force to the robot to affect its behaviors at any time, as shown in Fig. 2. We design a kinesthetic correction hardware setup with a detachable handle for human to hold when correcting, and allows easy tool-swapping for different tasks. We run a compliance controller (specifically admittance control) in the background to respond to both contact forces and human correction forces, allowing the human to influence but not completely override the policy execution. The admittance controller forms a virtual spring-mass-damper system around the base policy output $q_{ref}$:
    $M\ddot{q}=K(q_{ref}-q)-D\dot{q}+F,$ (1)
    where $M,K$ and $D$ denotes the virtual inertia, stiffness and damping of the compliant system, $F$ represents the combined force feedback from both external contacts and human correction. The admittance controller uses a constant stiffness $\sim$$1000\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$ to allow easy human intervention and ensure accurate tracking.
@@
Figure 3: Illustration of the intention misinterpretation caused by tracking error. In both cases, the robot feedback (and thus the human correction motion) are exactly the same. However, the recorded correction motion direction is very different. (A) recorded correction is correct when the tracking error is small. (B) recorded correction motion has wrong direction when there is large tracking error caused by both control and latency.
Additionally, comparing with the conference version of this paper CRDagger2025, we identified and fixed an issue of intention misinterpretation: When the human correction motion is smaller than the robot tracking error, the human intention can be misinterpreted and the residual policy would learn the wrong correction motion. This is illustrated in Fig. 3. With our Compliant Intervention Interface, human correction motion is computed as the difference between robot position feedback and the base policy output. A subtle assumption here is that the robot position before intervention accurately reflects the base policy behavior, i.e. the tracking error should be negligible. When the tracking error is larger than the delta corrective motion from human intervention, the computed correction motion could be in the opposite direction of the human’s true intended correction direction. To avoid intention misinterpretation even when human correction is small, we implement two features on top of CRDagger2025 to reduce the tracking error: 1) we add a velocity tracking term in the compliance control law in addition to the position tracking term used in 1:
$M\ddot{q}=K(q_{ref}-q)+D(\dot{q}_{ref}-\dot{q})+F,$ (2)
where the velocity reference $\dot{q}_{ref}$ is computed via finite difference from the position reference $q_{ref}$. This reduces the tracking error spatially. 2) We added a look-ahead time to robot feedback when computing the correction motion $q_{delta}$:
$q_{delta}[t]=q[t]-q_{ref}[t-\Delta t],$ (3)
where $\Delta t$ is empirically set by the estimation of hardware latency. The look-ahead time reduces the tracking error temporally. Together, the largest tracking error reduces from $30\text{\,}\mathrm{m}\mathrm{m}$ to below $5\text{\,}\mathrm{m}\mathrm{m}$ during fastest robot motion in our experiments.
@@ 3.2 Compliant Residual Policy
Compliant residual policy formulation. Our policy directly learns corrective behavior from the human delta correction data, as shown in Fig. 4. It takes as input the same visual and proprioceptive feedback as the base policy but with a shorter horizon. It also takes in an extra force modality, which is available using our Compliant Intervention Interface. The policy outputs five frames of actions at a time, corresponding to $0.1\text{\,}\mathrm{s}$ of execution time when running at $50\text{\,}\mathrm{H}\mathrm{z}$. The action space is 15-dimensional: the first nine dimensions represent the SE3 delta pose from the base policy action to the robot pose command chi2023diffusion, while the later six dimensions represent the expected wrench (force and torque) the robot should feel from external contacts. Both the robot pose command and the expected wrench are sent to a standard admittance controller for execution with compliance.
The residual policy directly uses the base policy’s frozen image encoder radford2021learning; amir2021deep; xu2023jacobinerf to extract an image embedding, a temporal convolution network van2016wavenet to encode the force vectors, followed by fully-connected layers to decode actions.
@@
Training strategy. In prior work, a residual policy is trained either in simulation with RL ankile2024imitation; yuan2024policy to give it sufficient coverage of the state distribution, or in the real world with pre-collected behavior cloning data luo2024precise. In this work, we train the Compliant Residual Policy completely on the small amount of new real-world correction data with the following strategies:
@@
  • Prioritize correction data over no-correction (zero residual action) data. Similar to liu2022robot, we alter the data sample frequency during training based on whether they have human correction or not. Specifically, since the moment of correction start indicates where the current policy performs badly followed by immediate action to fix it, we sample data more frequently for a short period immediately after correction starts. Our real-world ablations (§ 4.5) demonstrate that our training strategies improve the quality of the residual policy and the overall success rate.
@@ 4.1 Contact-Rich Manipulation Tasks
This task is challenging as it involves rich use of physical contacts and forceful strategies hou2020manipulation. A position-only strategy always fails immediately by triggering large forces, so we execute all policies through the same admittance controller. The task success requires high precision in both motion and force to accurately align the fingers with the gap upon insertion, and to provide enough force to rotate heavy books and make the books stand firmly.
@@
Belt Assembly: As shown in Fig. 6 (a), this task is to assemble a thin belt onto two pulleys, which is part of the NIST board assembly challenge kimble2020benchmarking. Starting with the belt grasped by the gripper, the robot needs to first thread the belt over the small pulley, next move down while stretching the belt to thread its other side on the big pulley, then rotate $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ around the big pulley to tighten the belt, and finally pull up to release the belt from the gripper. The base policy is trained with 405 demonstrations.
@@
Figure 7: Cable Routing Task. (a) Policy rollout of [Compliant Residual] policy trained with [On-Policy Delta] data. (b) Typical failure case of the base policy: missing a clip.
Cable Routing The task includes routing a USB cable around three clips, as shown in Fig. 7 (a). The USB cable has one fixed end and a free end. The cable starts being grasped, and can slide easily through a groove on the finger as the hand moves. The base policy is trained with 200 demonstrations.
We make this task challenging by using cables with different appearances and stiffness than those seen by the base policy. A change in cable position or physical property can easily cause the cable to be too high and miss a clip (Fig. 8 (b)), or too low and get stuck by a clip. We additionally introduce difficulty by using clips that are visually different from the one used in base policy training, testing our system’s generalization capability. We ran 20 rollouts across five cables and four different fixed end locations.
Gear Insertion The task involves inserting a plastic gear onto a metal axis while mating it to a nearby gear, which is also a part of the NIST board assembly challenge kimble2020benchmarking. The base policy is trained with 150 demonstrations.
The task is challenging due to the tiny tolerance between the gear and the axis. It is also different than the previous tasks as the human correction motions often have small magnitudes. We introduce additional difficulties by elevating the base board by $3\text{\,}\mathrm{c}\mathrm{m}$. We ran 20 rollouts across five different base locations.
Figure 8: Gear Insertion Task. (a) Policy rollout of [Compliant Residual] policy trained with [On-Policy Delta] data. (b) Typical failure case of the base policy: misaligning the gear with the axis and failing to mate with the nearby gear.
@@ 4.2 Base Policy and its Failure Modes
We trained a diffusion policy chi2023diffusion that takes in past images from a wrist-mounted camera and robot proprioception observations, and predicts a future position-based action trajectory. To isolate the contribution of force inputs versus human corrections, we trained diffusion policies with and without force inputs as baselines for the belt assembly task.
@@ 4.3 Baselines
  • Take-over-Correction (HG-DAgger) kelly2019hg is another common correction strategy where human demonstrators monitor policy execution and take complete control when failures are anticipated. We implement Take-over-Correction on our Compliant Intervention Interface by cleaning up command buffer to the compliance controller and switching stiffness to zero upon correction starts, so the robot policy does not affect the robot during correction. However, as explained in § 3.1, take-over correction introduces an abrupt transition around control authority switching, which may cause distributional discontinuities in the training data.
@@
  • Fintune Policy with KL Regularization: A recent method fan2023dpok that stabilizes finetuning training by encouraging the predicted action to be close to the training data distribution.